# Bridging Earth and Space: A Survey on HAPS for Non-Terrestrial Networks

### v1

@@ V-H Interference Management and Beamforming
In [171], a space-division multiple-access scheme was proposed with a multi-beam antenna and beamforming algorithm that adapts array weight vectors to improve performance and reduce interference. An improved elliptical cell design balanced SNR by broadening inner beams and narrowing outer ones, expanding inner cell coverage while shrinking outer areas. This reduced the number of required cells while maintaining coverage quality. The integration of space-time block coding with multi-beamforming further enhanced coding and diversity gains without extra bandwidth.
In [188], a dual-polarized shared aperture antenna for Ku- and Ka-bands was proposed, developed through a four-step derivation process using slot aperture coupling to achieve broadside radiation in both bands. Based on array factor theory, the optimal inter-element spacing was determined, resolving the trade-off between beam steering angles and array spacing in dual-band systems.
In [189], RA for HAPS-to-terrestrial and HAPS-to-satellite uplinks was optimized by addressing interference from uplink back lobes affecting GUEs. Unlike traditional zero forcing, which allocates random nulls, the proposed method generates targeted nulls based on GUE locations using a codebook of antenna patterns tied to weight matrices, enabling dynamic adaptation to reduce interference and improve communication quality.
In [129], traffic-aware RA was proposed for integrated HAPS-terrestrial systems, addressing the inefficiency of static methods that ignore load variations. The scheme adapts RA through interference coordination based on traffic distribution and network topology, achieving higher throughput than fixed strategies.
In [190], uplink performance in a multi-HAPS system jointly covering the same cell area was studied. To address the uplink bottleneck, the work explored spectrum reuse and interference mitigation using space-time diversity techniques, including MIMO and SIMO receive diversity.
@@
In [191], a multi-cell HAPS system was analyzed, where capacity is limited by feeder link bandwidth. To enhance spectrum efficiency, a multi-gateway approach was proposed, allowing gateways to share feeder link frequencies to reduce interference and increase capacity. Building on this, [192] introduced interference cancellation techniques for the multi-gateway feeder link, further improving transmission quality and overall communication performance.
To further reduce interference, the work in [193] introduced a null-sweeping technique where the HAPS dynamically adjusts null points, while terrestrial BSs schedule GUEs accordingly. This coordinated approach improved the SINR for GUEs without degrading the HAPS link quality, resulting in up to a 15% increase in ground system capacity.
V-I Energy Efficiency
@@
In [194], an adaptive link-state advertisement routing algorithm was proposed for integrated HAPS/satellite networks to support energy-efficient emergency communications. Designed for slow-flat Rician fading channels, the method estimates link information and selects the most energy-efficient path, reducing terminal power consumption while preserving connectivity.
In order to improve energy efficiency, the authors in [195] proposed a low-power-consumption decoding and forward relay system via HAPS. The results demonstrated that, in certain channel conditions, a relay-based system can outperform a non-relay-based system in terms of energy efficiency.
In [73], energy efficiency of a HAPS-based RAN was evaluated against a terrestrial RAN using a numerical model based on real traffic trends. Results showed that leveraging HAPS features—self-sustainability, high altitude, and wide coverage—can cut energy consumption by up to nearly 30%.
In [107], HAPS were studied as repeaters and aerial BSs, with energy efficiency assessed via consumption factor theory in single- and multi-cell scenarios using Monte Carlo simulations. Results showed both architectures achieve good downlink spectral efficiency, though coverage is constrained by GUE transmit power and antenna gain. Performance in uplink and downlink can be further improved with directed GUE antennas on LTE band 1.
In [196], a LEO satellite network using a solar-powered HAPS relay was studied with the goal of minimizing energy consumption for downlink data transmission. In the non-causal case, an energy-minimization problem was formulated and solved using future channel/energy knowledge, while in the causal case, a greedy online scheduling method was proposed, significantly improving throughput and energy efficiency.
In [197], an energy-optimized trajectory planning problem was formulated for high-altitude, long-endurance solar-powered aircraft. Using a 6-DoF dynamic model, the optimization maximized energy storage by combining battery and gravitational potential energy, solved via a nonlinear simplex search and interior point methods. The resulting optimal trajectories trained an Adaptive Neuro-Fuzzy Inference System for real-time planning, achieving a 15% gain in energy storage.
The study in [198] proposed a cost-effective power control algorithm for cognitive satellite-terrestrial networks with outdated CSI, where the satellite network opportunistically reuses the TN spectrum. This scheme was formulated as an energy efficiency maximization problem, subject to interference power constraints at the TN, and rate, energy and latency constraints at the satellites. The average transmit power was also determined by solving a nonlinear concave optimization problem.
In [199], a vertical HetNet with a HAPS as an SMBS, alongside an macrocell base station (MBS) and multiple small base stations, was studied to improve energy efficiency. An algorithm was proposed to minimize consumption by leveraging the HAPS and dynamically managing SBS sleep modes based on real-time traffic load, reducing power use by 30% while maintaining QoS. In [200], energy efficiency for data collection in a UAV network supported by a HAPS aerial BS was optimized. The approach dynamically clustered the target area, selected sink sensing devices, adjusted transmit power, and managed UAV operations. A centralized reinforcement learning (RL) framework based on proximal policy optimization (PPO) jointly optimized UAV trajectories and power, while an Affinity Propagation clustering algorithm identified optimal sink devices for device-to-UAV communication.
The research in [201] focused on energy-efficient RA in SAGINs for Internet of remote things (IoRT) using UAV relays. The goal was to improve energy efficiency by optimizing sub-channels selection and the uplink transmission power using Lagrangian dual decomposition, and UAVs deployment via sequential convex approximation. The proposed approach achieved more than 21.9% improvement in terms of energy efficiency.
The authors in [202] proposed a wireless backhaul solution using a RIS mounted on a HAPS to support UAV-BSs during sudden traffic surges in urban areas. To overcome signal blockages from urban obstacles, the aerial RIS reflects backhaul signals and adjusts the phase of each RIS element, position, and array partitioning, ensuring reliable connectivity. The use and optimization of RIS maximize energy efficiency and guarantees stable backhaul links for each UAV.
@@
The study in [203] explored the use of HAPS in SAGINs to support data transmission between LEO satellites and smart devices for IoRT applications. The proposed approach jointly optimized RA and HAPS deployment to minimize the power consumption. By breaking down the complex problem into two sub-problems, the authors found a near-optimal solution. Simulation results showed a significant reduction in power consumption compared to conventional RA strategies, while still satisfying smart device rate requirements.
While MEC enables GUEs to offload resource-intensive tasks, the stringent network requirements of 6G applications may exceed the capabilities of terrestrial MEC servers. To address this issue, the work in [204] proposed to use aerial MEC servers to support GUEs and emergency communications. The authors introduced a hierarchical deep federated learning RA scheme for intelligent aerial-enabled smart city cyber-physical system, aiming at optimizing energy consumption while preserving task offloading privacy and maintaining QoS in dynamic environments.
V-J Machine Learning
@@
RL is increasingly applied to NTNs to optimize resource allocation, platform positioning, and trajectory control. By leveraging Markov decision process (MDP) frameworks, RL enables aerial platforms like HAPS and UAVs to autonomously adapt and enhance network performance under real-world constraints. The survey [206] reviewed RL applications for NTNs—including LEO satellites, HAPS, and UAVs—focusing on resource allocation, platform positioning, and trajectory control under NTN-specific constraints. Using MDP frameworks, it provided a taxonomy of current research, assessed study realism, and identified open challenges to narrow the simulation-to-reality gap.
In this context, the authors in [207] considered an heterogeneous network with UAV LAPSs and HAPS, aiming at optimizing fairness and QoS for GUEs. The proposed approach combines DRL with fixed-point iteration to determine the optimal UAVs positions, 3D trajectories, and channel allocation strategies based on real-time traffic load. A RL-based approach is proposed to optimize UAV base station positioning in emergency scenarios where rapid deployment is critical [208]. By leveraging radio signal sensing and PPO, UAVs dynamically learn and generalize positioning strategies across diverse user mobility patterns. Their numerical evaluations demonstrate the algorithm’s adaptability, consistently achieving throughput above 7 Mbps in simpler scenarios — close to the maximum achievable rate of 8 Mbps — and maintaining robust performance even in more complex, dynamic environments such as multi-cluster and hotspot user movements.
The study in [209] highlighted the advantages of optimizing user scheduling in space-HAPS-ground networks by ensembling deep neural networks. In particular, it was demonstrated that, when the ensemble size increases, the suggested approach performs better than all existing offline optimization strategies, with a sum-rate improvement of up to 26%.
Considering RL techniques, the authors in [206] addressed key challenges such as limited resources, integration with TNs, and the need for intelligent control of aerial BSs. They also provided a detailed taxonomy of RL applications in NTN. The study [210] develops a deep RL-based offloading strategy for HAPS-assisted vehicular networks. The partial task offloading problem is formulated as a MDP and solved using the deep deterministic policy gradient (DDPG) algorithm. Among all evaluated schemes, the DDPG-based approach achieves the highest task success rate, with only a 0.7% failure rate when the task size is 1.2 Mbits. The authors in [175] focused on an integrated satellite-HAPS-TN architecture to enhance network performance by using RISs for signal reflection. Specifically, they optimized transmit beamforming and RIS phase shifts, a complex problem due to RIS constraints. The study employed a DRL algorithm for effective online optimization, demonstrating superior performance, speed, and feasibility in real-time decision-making through simulations.
In [140], a deep SARSA-based RL algorithm was proposed for user scheduling in a wireless network with a HAPS backup. Modeled as an MDP, it operated with outdated CSI to maximize sum-rate while minimizing active antennas at the HAPS for energy savings. Results showed near-optimal sum-rate performance, and a heuristic method was also introduced for solving the scheduling problem.
In [211], a dynamic coverage path planning algorithm for multi-airship formations was proposed to overcome the limitations of static approaches for mobile platforms such as fixed-wing aircraft and LEO satellites. Using RL, the method stored experiences in a pool for rapid updates via semi-centralized exploration and centralized playback, with task-specific rewards improving network coverage.
In [212], a multi-agent deep Q-learning algorithm was proposed for HAPS power control to reduce outage probability while satisfying interference constraints with TNs. A double deep Q-learning variant was also introduced to avoid action-value overestimation. Simulations showed that agents collaboratively learned a near-optimal power control policy through effective state, reward, and training design.
In [213], an RA method for uplink HAPS-LEO communication using multi-carrier NOMA was proposed. The problem was formulated as a mixed-integer nonlinear program under power and channel constraints, and solved with a multi-agent double deep Q-network employing prioritized experience replay and priority sampling to handle dynamic resources and short satellite visibility.
In [214], HAPS were studied as flying BSs serving GUEs with NOMA in a downlink THz system. The method optimized power allocation, bandwidth, antenna beamwidth, and HAPS altitude to improve transmission rates while meeting data requirements. A DRL-based DDPG algorithm was applied to manage network dynamics and complexity, maximizing long-term rewards with continuous decision variables.
Recent work shows that federated learning (FL) can strengthen HAPS-enabled networks by enhancing privacy and lowering communication overhead. Approaches like support vector machine (SVM)-based FL for task offloading and hierarchical FL with HAPS as distributed servers improve model accuracy, energy efficiency, and reduce training time for large-scale applications.

### v2

@@ VI-H Interference Management and Beamforming
Later, Guan et al. [82] introduced a space-division multiple-access scheme with a multi-beam antenna and beamforming algorithm that adapts array weight vectors to improve performance and reduce interference. An improved elliptical cell design balanced SNR by broadening inner beams and narrowing outer ones, expanding inner cell coverage while shrinking outer areas. This reduced the number of required cells while maintaining coverage quality. The integration of space-time block coding with multi-beamforming further enhanced coding and diversity gains without extra bandwidth.
@@
Fujii, Ohta and Fujii [70] analyzed a multi-cell HAPS system, where capacity is limited by feeder link bandwidth. To enhance spectrum efficiency, a multi-gateway approach was proposed, allowing gateways to share feeder link frequencies to reduce interference and increase capacity. Building on this, the authors [69] introduced interference cancellation techniques for the multi-gateway feeder link, further improving transmission quality and overall communication performance. To further reduce interference, Ishikawa et al. [102] introduced a null-sweeping technique where the HAPS dynamically adjusts null points, while terrestrial BSs schedule GUEs accordingly. This coordinated approach improved the SINR for GUEs without degrading the HAPS link quality, resulting in up to a 15% increase in ground system capacity.
Considering HAPS-aided SAGIN architecture, Kawamoto et al. [120] aimed to mitigate the interference between HAPS-to-GUEs links and back lobes from HAPS-to-satellite links. Unlike traditional zero forcing, which allocates random nulls, the proposed method generates targeted nulls based on GUE locations using a codebook of antenna patterns tied to weight matrices, enabling dynamic adaptation to reduce interference and improve communication quality.
VI-I Energy Efficiency
@@
A key challenge for emergency communication is ensuring energy-efficient transmissions due to the limited power resources available during large-scale disasters. Integrated HAPS/satellite networks can help overcome this issue by optimizing power consumption while maintaining robust connectivity. In this context, Dong et al. [62] proposed an adaptive link-state advertisement routing algorithm designed for a slow-flat Rician fading channel to estimate link information. This approach can reduce energy consumption at the terminal by selecting the most energy-efficient path.
In order to improve energy efficiency, Rallage and Kandeepan [177] proposed a low-power-consumption decoding and forward relay system via HAPS.The results demonstrated that, in certain channel conditions, a relay-based system can outperform a non-relay-based system in terms of energy efficiency. Prior to that, Xing et al. [232] studied HAPS as repeaters and aerial BSs, with energy efficiency assessed via consumption factor theory in single- and multi-cell scenarios using Monte Carlo simulations. Results showed both architectures achieve good downlink spectral efficiency, though coverage is constrained by GUE transmit power and antenna gain. Performance in uplink and downlink can be further improved with directed GUE antennas on LTE band 1. Also, Alfattani et al. [22] proposed an energy-optimization algorithm for HAPS, enabling mode switching to reduce consumption and extend loitering duration. The modes include a low-power HAPS-RIS mode for passive communication and a HAPS-SMBS mode for enhanced computing, caching, and communication. The system prioritizes the passive mode, activating higher-energy modes only when additional capacity is required. Finally, Salamatmoghadasi, Mehrabian, and Yanikomeroglu [185] studied a vertical HetNet with a HAPS as an SMBS, alongside an macrocell base station (MBS) and multiple small base stations, to improve energy efficiency. An algorithm was proposed to minimize consumption by leveraging the HAPS and dynamically managing SBS sleep modes based on real-time traffic load, reducing power use by 30% while maintaining QoS.
@@
Li et al. [135] studied energy-efficient RA in SAGINs for remote IoT using UAV relays. The goal was to improve energy efficiency by optimizing sub-channels selection and the uplink transmission power using Lagrangian dual decomposition, and UAVs deployment via sequential convex approximation. The proposed approach achieved more than 21.9% improvement in terms of energy efficiency. Later, Zhu et al. [253] explored the use of HAPS in SAGINs to support data transmission between LEO satellites and smart devices for remote IoT applications. The proposed approach jointly optimized RA and HAPS deployment to minimize the power consumption. By breaking down the complex problem into two sub-problems, the authors found a near-optimal solution. Simulation results showed a significant reduction in power consumption compared to conventional RA strategies, while still satisfying smart device rate requirements. Finally, Zhao and Chang [247] optimized energy efficiency for data collection in a UAV network supported by a HAPS aerial BS. The approach dynamically clustered the target area, selected sink sensing devices, adjusted transmit power, and managed UAV operations. A centralized reinforcement learning (RL) framework based on proximal policy optimization (PPO) jointly optimized UAV trajectories and power, while an Affinity Propagation clustering algorithm identified optimal sink devices for device-to-UAV communication.
Recently, Song et al. [201] evaluated the energy efficiency of a HAPS-based RAN against a terrestrial RAN using a numerical model based on real traffic trends and showed that leveraging HAPS features—self-sustainability, high altitude, and wide coverage—cuts energy consumption by up to 30%.
VI-J Machine Learning
@@
RL is increasingly applied to NTNs to optimize resource allocation, platform positioning, and trajectory control. By leveraging Markov decision process (MDP) frameworks, RL enables aerial platforms like HAPS and UAVs to autonomously adapt and enhance network performance under real-world constraints. In their survey, Naous et al. [163] reviewed RL applications for NTNs—including LEO satellites, HAPS, and UAVs—focusing on resource allocation, platform positioning, and trajectory control under NTN-specific constraints. Using MDP frameworks, it provided a taxonomy of current research, assessed study realism, and identified open challenges to narrow the simulation-to-reality gap.
In this context, Nguyen et al. [165] developed a deep RL-based offloading strategy for HAPS-assisted vehicular networks. The partial task offloading problem was formulated as a MDP and solved using the deep deterministic policy gradient (DDPG) algorithm. Among all evaluated schemes, the DDPG-based approach achieves the highest task success rate, with only a 0.7% failure rate when the task size is 1.2 Mbits. Dahrouj, Liu, and Alouini [57] highlighted the advantages of optimizing user scheduling in space-HAPS-ground networks by ensembling deep neural networks. In particular, it was demonstrated that, when the ensemble size increases, the suggested approach performs better than all existing offline optimization strategies, with a sum-rate improvement of up to 26%. Also, Sharifi et al. [195] proposed a deep SARSA-based RL algorithm for user scheduling in a wireless network with a HAPS backup. Modeled as an MDP, it operated with outdated CSI to maximize sum-rate while minimizing active antennas at the HAPS for energy savings. Results showed near-optimal sum-rate performance, and a heuristic method was also introduced for solving the scheduling problem. Jo et al. [111] introduced a multi-agent deep Q-learning algorithm for HAPS power control to reduce outage probability while satisfying interference constraints with TNs. A double deep Q-learning variant was also introduced to avoid action-value overestimation. Simulations showed that agents collaboratively learned a near-optimal power control policy through effective state, reward, and training design. Later, a dynamic coverage path planning algorithm was proposed by Liu et al. [141] for multi-airship formations to overcome the limitations of static approaches for stratospheric airship formations in dynamic coverage scenarios. Using RL, the method stored experiences in a pool for rapid updates via semi-centralized exploration and centralized playback, with task-specific rewards improving network coverage. Meanwhile, Arani, Hu, and Zhu [28] considered a heterogeneous network with UAVs and HAPS as aerial BSs, aiming at optimizing fairness and QoS for GUEs. The proposed approach combines DRL with fixed-point iteration to determine the optimal UAVs positions, 3D trajectories, and channel allocation strategies based on real-time traffic load. Minani et al. [157] introduced an RA method for uplink HAPS-LEO communication using multi-carrier NOMA. The problem was formulated as a mixed-integer nonlinear program under power and channel constraints, and solved with a multi-agent double deep Q-network employing prioritized experience replay and priority sampling to handle dynamic resources and short satellite visibility. Later, Le et al. [130] studied HAPS as flying BSs serving GUEs with NOMA in a downlink THz system. The method optimized power allocation, bandwidth, antenna beamwidth, and HAPS altitude to improve transmission rates while meeting data requirements. A DRL-based DDPG algorithm was applied to manage network dynamics and complexity, maximizing long-term rewards with continuous decision variables.